if c=20 +0.9Y c=consumption y=disposable income=800 billion what is the multiplier?

Question 2 If C = 20 + 0.99 C= Consumption Y = disposable income = $800 billion What is the Multipler? Ob.0.1

also calculate MPS MPC APS APC
and what is the value of saving for the nation?
Answer #1

Ans: c)10

Multiplier = 1/(1-MPC)

where, MPC = Marginal propensity to consume = coefficient of Y in consumption function = 0.9

Therefore, multiplier = 1/(1-0.9)

= 1/0.1

= 10